OpenStudy (anonymous):

If I know the lengths of two sides of a right triangle, how do I find the third? i have to answer it and do an example plzz help

OpenStudy (amistre64):

pythag thrm a^2 + b^2 = c^2

OpenStudy (apoorvk):

where c is the largest side, opposite to the right angle, and is called hypotenuse

OpenStudy (amistre64):

details smetails lol

OpenStudy (amistre64):

you ever seen euclids "proof"?

OpenStudy (amistre64):

 something like that :)

Directrix (directrix):

Pythagorean Theorem The square of the hypotenuse of a right triangle is equal to the sum of the squares of the other two sides. If you are given 3 and 4 as the legs of a right triangle and want to find the length of the hypotenuse, then use the Pythagorean Theorem. h^2 = 3^2 + 4^2 h^2 = 9 + 16 h^2 = 25 h = 5 Note: -5 is another solution to h^2 = 25 but -5 cannot be the length of a side of a triangle.